The parametric equations for a curve in the x-y plane are x=2+t^2 and y=4-3t. Determine the points where the curve intersects the x-axis.
Thank you!

Answer by stanbon(57307)

The parametric equations for a curve in the x-y plane are x=2+t^2 and y=4-3t. Determine the points where the curve intersects the x-axis.
-----
When the curve hits the x-axis y = 0.
---
Solve: 4-3t = 0
t = (4/3)
----
Then x = 2+t^2
x = 2+(4/3)^2
x = 18/9 + 16/9
x = 34/9
-------------------------
The curve intersects at (34/9,0)
==================================
Cheers,
Stan H.